# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 3.5.1 Basic idea
However, if we merge the first 3 states in group 1 into 1 new state, and keep the 4th state in group 2 as it is, then we get a coarse-grained Markov model with 2 states, with TPM:
@@
Now the future states of the new system can be fully determined by the current states. It shows that we can eliminate the uncertainty of an undeterministic system by doing coarse-graining over the system states.
3.5.2 About Coarse-graining
The example above illustrates how a system’s determinism can increase as it is coarse-grained. Coarse-graining is a process that simplifies the description of a system by grouping its components into larger, more slowly varying units. It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of microscale interactions. Unlike dimension reduction techniques, such as PCA and SVD, coarse-graining takes into account the system’s features at different spatial and temporal scales. Coarse-graining also differs from renormalization. In physics, the method of renormalization group is invented to avoid the infinity in integrals, it is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro- levels Christensen and Moloney (2005); McComb (2003). However, coarse-graining methohd does not have this requirement in general. Although both techniques are targeting to describe a system from a coarse-grained level, however, coarse-graining are more focused on the states of the system, while renormalization cares of the dynamics, the system rules or partition function etc, which is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking. These are aspects that coarse-graining typically does not consider.
Given a transitional probability matrix, there are a large amount of ways to partition the state space in order to coarse-grain the system. What kind of coarse-graining strategies are more reasonable? In Hoel et al. (2013) and Hoel (2017), the authors suggest to select an appropriate coarse-graining method by maximization effective information of the coarsed TPM. In history, there is another criterion to select coarse-graining strategy called lumpability. Where lumpability of a TPM refers to the coarse-grained TPM exhibits similar dynamics and observation statistics Kemeny and Snell (1983). Commutativity is another requirement which needs the operations of coarse-graining and state transition being commutative Pfante et al. (2014). There are also other discussions or reasonable model reduction of Markov or hidden Markov models Kotsalis et al. (2008); White et al. (2000); Wolpert et al. (2014)
3.5.3 About Effective Information
In order to quantify emergence, a widely used method is to calculate the mutual information between the past states and future states of a system. This mutual information defines the upper limit of the information being transferred from past to future. When the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
One limitation of utilizing mutual information is that the resulting value may fluctuate in response to changes in the joint probability of data. This can make it difficult to obtain a consistent outcome when the system is exposed to different inputs.
To address this limitation, Hoel introduced the concept of effective information. This concept measures the causal effect of the current state of a system on its future state, also based on mutual information. Effective information is a scoring metric that is calculated from the system’s transition probability matrix (TPM). The TPM does not change solely based on input data.
Assuming that the current state of the system is in the maximum entropy distribution $I_{D}$ (uniform distribution), the TPM predicts that the state distribution of the system at the next moment (i.e., in the future) will be $E_{D}$. The effective information $EI$ of the system is then defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
$EI=I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{X_{t}\sim U}$ (7)
$X_{t}\sim U$ represents the probability of the random variable $X_{t}$ following a uniform distribution. It can be readily demonstrated that the effective information ($EI$) can be fully described by the transition probability matrix ($TPM$) (see Appendix 7).

### 25 Feb 2024 v3

@@ 3.5.1 Basic Idea
However, if we merge the first three states in group 1 into one new state and keep the fourth state in group 2 as it is, we obtain a coarse-grained Markov model with two states, represented by the following TPM:
$S_{M}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ (5)
Now, the future states of the new system can be fully determined by the current states. This shows that we can eliminate the uncertainty of a nondeterministic system by performing coarse-graining over the system states.
3.5.2 About Coarse-Graining
The example above illustrates how a system’s determinism can increase as it is coarse-grained. Coarse-graining is a process that simplifies the description of a system by grouping its components into larger, more slowly varying units. It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of micro-scale interactions. Unlike dimension reduction techniques, such as PCA and SVD, coarse-graining takes into account the system’s features at different spatial and temporal scales. Coarse-graining also differs from renormalization. In physics, the renormalization group method was invented to eliminate infinities in integrals. It is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro-levels Christensen and Moloney (2005); McComb (2003). However, the coarse-graining method does not have this requirement in general. Although both techniques are designed to describe a system from a coarse-grained level, coarse-graining focuses more on the states of the system, whereas renormalization cares about dynamics, system rules, partition functions, etc. Renormalization is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking. These are aspects that coarse-graining typically does not consider.
Given a transitional probability matrix (TPM), there are many ways to partition the state space for coarse-graining the system. What kind of coarse-graining strategies are more reasonable? In Hoel et al. (2013); Hoel (2017), the authors suggested selecting an appropriate coarse-graining method by maximizing the effective information of the coarse TPM. In the literature, there is another criterion for selecting a coarse-graining strategy called lumpability, where the lumpability of a TPM refers to the coarse-grained TPM exhibiting similar dynamics and observation statistics Kemeny and Snell (1983). Commutativity is another requirement that requires the operations of coarse-graining and state transitions to be commutative Pfante et al. (2014). There are also other discussions or reasonable model reductions of Markov or hidden Markov models Kotsalis et al. (2008); White et al. (2000); Wolpert et al. (2014).
3.5.3 About Effective Information
In order to quantify emergence, a widely used method is to calculate the mutual information between the past states and future states of a system. This mutual information defines the upper limit of the information being transferred from the past to the future. If the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
One limitation of utilizing mutual information is that the resulting value may fluctuate in response to changes in the joint probability of data. This can make it difficult to obtain a consistent outcome if the system is exposed to different inputs.
Effective Information (EI)
To address this limitation, Hoel introduced the concept of effective information to measure the causal effect of the current state on the future state of a system. Effective information is a scoring metric based on mutual information and can be calculated from the system’s transition probability matrix (TPM), which is invariant to the input data.
By “intervening” in the current state distribution to follow the uniform distribution (the distribution of maximum entropy), denoted as $I_{D}$, the TPM enables the prediction of the future state distribution $E_{D}$ at the next moment. The effective information ($EI$) of the system is defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
$EI\equiv I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{do(X_{t})\sim U}$ (7)
The probability distribution of the random variable $X_{t}$, representing the initial state, is denoted as $do(X_{t})\sim U$, indicating that it follows a uniform distribution. Hoel adopted the notion of the “do” operator from Judea Pearl’s causal analysis framework Hoel (2017). It is worth noting that in Pearl’s context, the do operator is typically used to assign a specific value to the intervened variable rather than applying a distribution Pearl (2009).
Another point that needs clarification is that the “do” operator used here is purely a mathematical construct that specifies the distribution of $X_{t}$. It does not imply the need for actual intervention in the system. However, the imaginary “do” operation is equivalent to a real intervention in the context of this scenario, given that the dynamical mechanism (TPM) is provided. Consequently, we can perform any desired intervention on the system, just as in computer simulations.
The second issue pertains to why we apply the “do” operator to a uniform distribution. In Hoel et al. (2013), Hoel et al. claimed that the distribution should be the maximum entropy distribution, which is the most reasonable selection for the input variable $X_{t}$ if we have no prior information about the input variable Jaynes (1957). As we know, uniform distribution can be derived by maximizing entropy if there is no constraint.
The authors of this review believe that applying the “do” operator to a uniform distribution ensures that the objective measured by EI solely reflects the dynamical mechanism itself, i.e., the TPM, and is independent of any input data.
This point can be clearer if we re-express EI as a function of the TPM of the system (see Appendix A for the detailed derivation):
$EI=\frac{1}{N}\sum_{i,j}TPM(i,j)\log_{2}\Big{(}\frac{N\times TPM(i,j)}{\sum_{k}TPM(k,j)}\Big{)},$ (8)
where $TPM(i,j)$ represents the transitional probability of the system from the state $i$ to the state $j$.
Therefore, effective information offers a solution to the aforementioned limitation of mutual information. Given that the TPM can capture the inherent nature of a system, EI is likewise an intrinsic property of the system.
However, if we apply the “do” operator to the system using different distributions, EI will depend on the chosen distribution, and certain transition probabilities of specific rows may carry greater weight in the average calculation. Consequently, deriving a simple expression solely based on the transitional probability matrix (TPM) may not be possible. Furthermore, if there is no “do” operator, the effects of the dynamical mechanism and the input distribution in EI would not be distinguishable or separated, as the mutual information is a function of the input distribution $p(X_{t})$.
The Derived Measures of EI
Furthermore, we can calculate the EIs for the TPMs of both macro- and micro-dynamics, and their difference is defined as the measure of causal emergence, that is,
$CE=EI(TPM_{M})-EI(TPM_{m}),$ (9)
where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the TPM of micro-level dynamics. $CE$ measures the degree of causal emergence. If $CE>0$, then causal emergence occurs; otherwise, it does not occur.
However, there is a limitation for EI and CE when comparing two dynamics that differ significantly in size. This is because the value of effective information relies on the number of possible states within the system, denoted as $N$, with an upper bound of $\log_{2}(N)$. To facilitate comparisons between different coarse-graining strategies and scales, effective information is often normalized, resulting in a metric known as the effect coefficient $Eff$.
$Eff=\frac{EI}{\log_{2}(N)}$ (10)
The value of the effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states. If the information is fully transferred, the effect coefficient is $1$.
In addition to characterizing causal emergence, the effect coefficient can be further broken down into two meaningful components: “$Determinism$”, which represents the certainty of a current state evolving into a certain state or diverging into multiple states in the next moment, and “$Degeneracy$”, which represents the possibility of multiple current states converging into one state in the next moment.
$Eff=Determinism-Degeneracy$ (11)
where both “$Determinism$” and “$Degeneracy$” can be defined in terms of the TPM (refer to Equations (8) and (10)):
$Determinism$ $=\frac{1}{N\log_{2}(N)}{\sum_{i,j}{TPM(i,j)\log_{2}(N\times TPM(i,j))}}$ (12)
$Degeneracy$ $=\frac{1}{N\log_{2}(N)}{\sum_{i,j}TPM(i,j)\log_{2}(\sum_{k}{TPM(k,j)})}$
In Figure 6, the square cells represent the elements of the TPM, and the grayscale areas represent the values of the TPM elements. Example (a) is a bijective system, meaning that all information from current states can be transferred to future states without loss. It is fully deterministic with zero degeneracy. Example (b) is an extreme case where all current states lead to only one future state, illustrating that high determinism does not necessarily imply a high effect coefficient. Systems (c) and (d) differ but have the same effect coefficients. Finally, system (e) is a coarse-grained version of either system (c) or (d), demonstrating two important points: first, different microscopic systems can be coarse-grained to the same macroscopic system, and second, causal emergence can be quantitatively captured from coarse-graining by the increased effect coefficient.
One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation (8). Therefore, Varley (2022) proposed the concept of flickering emergence, which decomposes EI into each term added in Equation (8). This new feature can characterize the local properties of Markov dynamics.
Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graining can change them. (a) Deterministic system with zero degeneracy; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from (c) or (d), resulting in a fully deterministic system with a high effect coefficient.
Comparison with Other Measures of Causation
Hoel’s theory of causal emergence is based on effective information, however, is the selection of EI necessary for measuring causation? Is causal emergence a phenomenon that depends on the selection of the measure of causation? To address this problem, in Comolatti and Hoel’s work Comolatti and Hoel (2022), a systematic comparison was conducted between effective information (EI) and other measures of causation that are widely applied in various fields, ranging from philosophy to genetics.
The findings of this comparison revealed that causal emergence is not merely a peculiar phenomenon limited to a specific measure. Instead, it exhibits commonalities and shared characteristics across different measures and fields of study. EI is not the sole measure for capturing causal emergence; there are other measures of causation that can also reveal the phenomenon of causal emergence. This suggests that the concept of causal emergence has broader applicability and relevance. We introduce this work in detail below.
Firstly, Hoel highlighted that causation is not merely a singular relationship between a cause and an effect. Instead, it encompasses two fundamental dimensions known as causal primitives: sufficiency and necessity.
The sufficiency aspect of causation refers to the scenario where the occurrence of the cause guarantees the occurrence of the effect. In other words, whenever the cause happens, the effect is also observed. This sufficiency dimension, denoted as $suff$, can be formally defined as the probability of the effect $e$ occurring given the condition that the cause $c$ has occurred:
$suff(e,c)=P(e|c).$ (13)
In addition, a necessary ($nec$) relation in causation refers to the absence of the effect implying the absence of the cause. In other words, when the effect does not occur, it indicates that the cause also does not occur. This can be understood as a kind of causal effect measure for the counterfactual. This necessary dimension, denoted as $nec$, can be quantitatively defined as the probability that the effect $e$ does not occur when the condition that the cause $c$ has not occurred is given by:
$nec(e,c)=1-P(e|C/c),$ (14)
where $C/c$ represents the set of all possible causes in $C$ but with the particular cause $c$ being excluded. Therefore, $P(e|C/c)$ represents the probability that $e$ occurs if other causes except $c$ occur.
With these causal primitives defined, Hoel compared different measures of causation, including Hume’s constant conjunction Hume (2016), Cheng’s causal attribution Cheng and Novick (1991), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009). The finding was that all of these measures can be expressed as the two causal primitives.
For example, Cheng’s causal attribution can be expressed as:
$CS_{cheng}=\frac{suff(e,c)+nec(e,c)-1}{nec(e,c)}$ (15)
With this understanding, a natural question arises: Can effective information (EI) be expressed using causal primitives? The answer is affirmative. However, to clarify this point, two important distinctions need to be made. Firstly, in most measures of causation, the causal variables are binary, whereas EI is defined for variables with multiple values. Secondly, EI is an information-theoretic measure, whereas others are probabilistic measures. Despite these distinctions, EI can still be expressed using causal primitives.
To understand this, let us examine the equivalent measure of EI: normalized effective information ($Eff$) in Equation (11). This measure contains two terms. The first term is $determinism$, and the second one is $degeneracy$. These two terms can be expressed as $suff$ and $nec$:
$determinism=1-\frac{\sum_{c\in C}P(c)H(e|c)}{\log_{2}N},$ (16)
where $H(e|c)\equiv-\sum_{e\in E}P(e|c)\log_{2}P(e|c)$ is the Shannon entropy of the conditional probability $P(e|c)$, which is $suff(e,c)$, and $P(c),\forall c\in C$ is the distribution of all causes. In the definition of EI (Equation 7), this distribution is intervened as a uniform distribution such that equal weights are assigned to causes $c\in C$. It is not hard to see that $determinism$ is an information metric for $suff(e,c)$ and is averaged for all causes. Furthermore, as $suff(e,c)=P(e|c)$ increases and approaches one, indicating the emergence of causal effect, the value of $H$ decreases, whereas determinism concurrently increases. Thus, the determinism term in EI plays a similar role to that of $suff$ in causal primitives.
Another term is $degeneracy$, which can be re-written as:
$degeneracy=1-\frac{H(e|C)}{\log_{2}N},$ (17)
where $H(e|C)\equiv-\sum_{e\in E}P(e|C)\log_{2}P(e|C)$ is also the Shannon entropy of the conditional probability $P(e|C)$, which is calculated by averaging the causal effect for all elements in $C$ as $P(e|C)=\sum_{c\in C}P(c)P(e|c)$. The entropy $H(e|C)$ serves as a measure of the average causal effect for counterfactuals, as $P(e|C)$ can be interpreted as $P(e|C/c)$. This is due to the approximation $P(e|C)\approx P(e|C/c)$ when the number of elements in $C$ is significantly large, and $C$ certainly encompasses $C/c$. Consequently, $degeneracy$ in $EI$ acts as the counterpart to $nec$ in causal primitives.
Therefore, EI or $eff$ is a valuable measure of causation, particularly in cases where the cause and effect variables are not limited to binary values.
Comolatti and Hoel Comolatti and Hoel (2022) conducted additional experiments to investigate the phenomenon of causal emergence using various Markovian dynamics, employing different measures of causation, as discussed above. Their findings revealed the widespread occurrence of causal emergence, regardless of the specific measure of causation employed.
3.5.4 Examples of Causal Emergence
The example above demonstrates a coarse-graining strategy that is intuitive, as it involves aggregation at the state level. However, in reality, coarse-graining often occurs among variables, and the combination of these variables produces different states. Let us look at a more complicated example of a boolean network, given in Hoel et al. (2013), and inspect the causal emergence when coarse-graining its state transition mechanism. A, B, C, and D are four binary variables whose state transition relations are shown in Figure 7a.
Figure 7: The mechanism and transition probability matrix of a boolean network. Each variable’s next state is determined by the combination of the current states of two other variables, as shown in the boolean network (a). If we examine an atomic subsystem with two inputs and one output, such as AB →C, the next state of C depends on the present states of A and B. If A = 0 and B = 0, the probability of C being 1 is 70%, whereas the probability of it being 0 is 30%, as indicated in the table (b). This same microscopic dynamics applies to other atomic subsystems such as AB →D, CD →A, and CD →B. When we expand to a subsystem of two inputs and two outputs, AB →CD and CD →AB, we obtain the transition probability matrix (TPM), as demonstrated in the table (c). Finally, we extend to the dynamics of the entire system ABCD →ABCD, whose TPM is represented as a heatmap (d). The effective information and relative metrics of the entire system are as follows: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
In order to demonstrate how different coarse-graining strategies can impact effective information, we can combine two micro-states into a single macro-state. For example, we can define $\alpha=(A,B)$ and $\beta=(C,D)$ as coarse-grained variables at the macro-level, where their values can be either ‘off’ or ‘on’. Using various aggregation strategies, as illustrated in (a)/(b) and (d)/(e) in Figure 8, we can construct the TPMs for the resulting coarse-grained systems. The TPM and the corresponding metrics are presented in Figure 8c,f, respectively.
The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.18$, respectively. It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), whereas it has decreased in the second case ($EI_{ma2}(0.18)<EI_{mi}(1.15)$). Thus, causal emergence is observed in the first coarse-graining scenario. This implies that (1) proper coarse-graining can lead to causal emergence in a macroscopic system, and (2) not all coarse-graining approaches result in higher effective information. This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well-defined quantification method. This gives us the opportunity to identify causal emergence in a given system. This raises the following question: What coarse-graining strategy can generate the maximum effective information? We address this question in Section 4.
Figure 8: An example showing how different coarse-graining strategies applied to the same boolean network, shown in Figure 7, can lead to different boolean dynamics. (a) Coarse-graining strategy 1. (b) Macro-dynamics resulting from strategy 1. (c) The corresponding heatmap of the TPM and the effective information metrics of the macroscopic system using method 1. (d) Coarse-graining strategy 2. (e,f) The macro-dynamics and TPM corresponding to coarse-graining strategy 2 in (d), respectively.
3.5.5 Extensions to Continuous Systems
Effective information and its associated metrics have been shown to be useful in quantifying causal emergence. However, there are some limitations to be aware of. First, the method is only applicable to discrete-state systems. Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate the effective information metrics. Efforts have been made to apply or extend the concepts of effective information to continuous systems Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
Ordinal Partition Network
One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones using the ordinal partition network (OPN) method Varley (2020). The basic idea is to transform the time-series data generated by a continuous dynamical system into a sub-time series of dimensionality $d$ with discrete values and then establish a transition probability matrix (TPM) between these discrete series. This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub-time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series. By employing rank orders, this method ensures that the values are restricted to integers within a certain range.
In Varley (2020), the author applied the OPN method to the Rossler chaotic attractor system. By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) was sensitive to the critical phase transition between period and chaos Varley (2020).
Causal Geometry
Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the concept of causal geometry Chvykov and Hoel (2020).
To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is,
$y=f(x)+\varepsilon,$ (18)
where $x$ is the input variable defined on the interval $[-L/2,L/2]$, $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a Gaussian noise with a zero mean and standard deviation $\epsilon\ll L$. Then, an approximate form of the effective information for such a continuous system can be obtained using the Gaussian integral.
The reason we set the domain of $x$ to $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented and to simplify the definition of a uniform distribution on it.
This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ with $x(t)$ and $x(t+1)$, respectively. The formula can also be extended to higher dimensions, with $\mathbf{x}$ and $\mathbf{y}$ represented in bold. Suppose $\mathbf{x}$ $\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}$ $\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers. Equation (19) can be generalized to the following form:
where $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, ||$\cdot$|| ⋅ | is the absolute value operation, and $\det$ is the determinant.
Note that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is, $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$. Thus, the term in the expectation in Equation (20) can be written as:
and this corresponds to the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
which measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ to changes in $\mathbf{x}$, where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$. Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, encompassing all possible distributions of $p(\mathbf{y}|\mathbf{x})$. This is the origin of the term “geometry” in this framework.
Finally, we can obtain an expression for EI with the Fisher information metric:
This formula can be generalized to cases where $p(\mathbf{y}|\mathbf{x})$ is a non-Gaussian distribution. Once the distribution function is known, we can obtain its Fisher information metric, and then EI can be calculated using Equation (23). The reason behind this is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
In Chvykov and Hoel (2020), the authors considered a more general case in which intervention noise is added to the input (intervention) variable $x$ in Equation (19). The intervention noise is denoted as $\xi\sim\mathcal{N}(0,\delta^{2})$, where $\delta$ is the standard deviation of $\xi$. In contrast, $\varepsilon$ is called observational noise distinguishing it from $\xi$, and is added to the output (observational) variable $y$ according to Equation (19). Finally, if two kinds of noise are both considered and if $L=1$ and $\epsilon\ll 1$, Equation (19) becomes:
@@ 3.5.5 Extensions to Continuous Systems
This is the formula for EI for the continuous mapping function (Equation (18)) given by Chvykov and Hoel (2020). Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation (18)) and a more discrete map with an adjustable parameter in the logistic function. They discovered that when the noise level is low, the continuous map can exhibit higher EI compared to the discrete map. However, as the noise level increases, discretizing the mapping function can lead to a model with higher EI. This phenomenon helps explain why digital circuits eventually outperform analog circuits in mitigating noise interference; the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we cannot control $\mathbf{y}$ by directly intervening $\mathbf{x}$. Instead, we can intervene $\mathbf{x}$ to influence $\theta$ and indirectly affect $\mathbf{y}$. Therefore, the three variables form a Markov chain: $\mathbf{x}\rightarrow\theta\rightarrow\mathbf{y}$.